A Hunter tries to hunt a monkey with a small Arrow with initial velocity u not the monkey is hanging on branch of tree of height H above the ground Hunter is l distance from the bottom of the tree monkey sees the arrow and immediately fall freely with zero initial velocity the the min speed of Arrow Hunter to succeed

Grade:12th pass

1 Answers

Eshan
5 years ago
Dear student,

For minimum speed of arrow, the arrow will hit the monkey at the bottom of the tree.
Hence the range of arrow=l
Displacement of monkey in time of flight of arrow=H
Hencel=\dfrac{2u^2sin\theta cos\theta}{g}
andH=\dfrac{1}{2}g(2\dfrac{usin\theta}{g})^2=\dfrac{2u^2sin^2\theta}{g}
Solving the two equations can give the value ofu